A 70 lb bag of rice is being pulled by a person by applying a force F at an angle θ as shown. The force required to drag the bag is given by: F(θ)=70μ/μsinθ+cosθ where μ=0.35 is the friction coefficient. (a) Determine F(θ) for θ= 5°, 10°, 15°, 20°, 25°, 30°, 35° (b) Determine the angle θ where F is minimum. Do it by creating a vector θ with elements ranging from 5°to 35°and spacing of 0.01. CalculateF for each value of θ and then find the maximum F and associated θ with MATLAB's built-in function max

1
Expert's answer
2022-02-17T13:36:50-0500

Solution.



"th=5:5:35; mu=0.35;\newline
% Part (a)\newline
a) \newline
F=70*mu./(mu*sind(th)+cosd(th)) \newline
% Part (b) \newline
b) \newline
thb=5:0.01:35; \newline
Fb=70*mu./(mu*sind(thb)+cosd(thb)); \newline
[Fm, i]=min(Fb);\newline
Fmin=Fb(i)\newline
atTheta=thb(i)\newline

Command \space Window:\newline
F =
 23.86\newline 
23.43 \newline
23.19 \newline
23.13 \newline
23.24 \newline
23.53 \newline
24.02\newline
Fmin =
 23.12\newline
atTheta =
 19.29 \newline
Answer: \newline
F =
 23.86\newline 
23.43 \newline
23.19 \newline
23.13 \newline
23.24 \newline
23.53 \newline
24.02\newline
Fmin =
 23.12\newline
atTheta =
 19.29"